OpenStudy (anonymous):

#1 Simplify 3a-5a+c-c #2 Solve 4-(2x-1)=x-13

OpenStudy (anonymous):

#1 Simplify 3a-5a+c-c = -2a #2 Solve 4-(2x-1)=x-13 4-2x+1=x-13 -2x+5=x-13 5+13=x+2x 18 = 3x 18/3=x x = 6

OpenStudy (gw2011):

1) 3a-5a+c-c=-2a 2) 4-(2x-1)=x-13 4-2x+1=x-13 5-2x=x-13 -2x-x=-13-5 -3x=-18 x=6